# Revisiting Pre-trained Language Models for Vulnerability Detection

### v1

@@
Abstract
The rapid advancement of pre-trained language models (PLMs) has demonstrated promising results for various code-related tasks. However, their effectiveness in detecting real-world vulnerabilities remains a critical challenge. While existing empirical studies evaluate PLMs for vulnerability detection (VD), their inadequate consideration in data preparation, evaluation setups, and experimental settings undermines the accuracy and comprehensiveness of evaluations. This paper introduces RevisitVD, an extensive evaluation of 17 PLMs spanning smaller code-specific PLMs and large-scale PLMs using newly constructed datasets. Specifically, we compare the performance of PLMs under both fine-tuning and prompt engineering, assess their effectiveness and generalizability across various training and testing settings, and analyze their robustness against code normalization, abstraction, and semantic-preserving transformations.
Our findings reveal that, for VD tasks, PLMs incorporating pre-training tasks designed to capture the syntactic and semantic patterns of code outperform both general-purpose PLMs and those solely pre-trained or fine-tuned on large code corpora. However, these models face notable challenges in real-world scenarios, such as difficulties in detecting vulnerabilities with complex dependencies, handling perturbations introduced by code normalization and abstraction, and identifying semantic-preserving vulnerable code transformations. Also, the truncation caused by the limited context windows of PLMs can lead to a non-negligible amount of labeling errors. This study underscores the importance of thorough evaluations of model performance in practical scenarios and outlines future directions to help enhance the effectiveness of PLMs for realistic VD applications.
Index Terms:
pre-trained language model, vulnerability detection, software security
I Introduction
@@
Although existing research has explored if PLMs can demonstrate strong performance in vulnerability detection (VD), limitations across various stages of the evaluation pipeline hinders an accurate reflection of PLMs’ capabilities: (i) Data Leakage. Most studies rely on evaluation datasets that inherently introduce data leakage, leading to biased estimations of the model performance in real-world scenarios. The randomly partitioning method used in existing studies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] often causes models to encounter duplicated code patterns between the training and test data. The temporal overlap between the pre-training data cut-off dates of PLMs and the commit dates of evaluation data is also frequently overlooked [16, 17, 11, 12, 13, 14, 15]. (ii) Limited Scope. The experimental setup and settings adopted by many studies are neither comprehensive nor aligned with real-world scenarios, resulting in unrepresentative conclusions. For example, some studies focus on models with constrained architectures and parameter scales [18, 1, 2, 3]. Also, models are often evaluated on samples with limited size, a narrow range of vulnerability types [19, 20, 21, 22, 23, 24, 11, 12, 13], or an unrealistically balanced distribution [25, 26, 27, 28, 21, 22, 23, 24]. Others focus exclusively on either fine-tuning [1, 2, 3, 4, 5, 6, 7, 8] or prompting engineering [25, 26, 27, 28, 21, 22, 23, 24], neglecting comparative analyses across different adaptation methods or considerations of the trade-offs between cost and performance of PLMs in VD. (iii) Superficial Evaluation. Existing studies [10, 16, 17, 11, 12, 13, 14, 15] primarily focus on performance comparisons, without thoroughly investigating how practical factors (e.g., code normalization, abstraction, transformation, truncation) influence the effectiveness of PLMs for VD. This creates a significant gap between performance estimates and real-world applications, failing to provide insights for enhancing PLMs’ true capabilities.
@@
Therefore, this paper Revisits the capabilities of PLMs for VD (RevisitVD) through an extensive and realistic evaluation that addresses the shortcomings of existing evaluation works. In particular, starting from data preparation, we discuss the limitations of existing work in selecting evaluation datasets, including insufficient consideration of data volume, the diversity of vulnerability types, and inherent labeling errors within the datasets. To address these issues, we introduce our reconstructed dataset, alongside a time-order-based dataset partitioning method to avoid data leakage caused by the random data partitioning commonly used in current VD research. Additionally, we highlight the risk that existing VD datasets may predate the cutoff date of PLMs’ pre-training data, potentially leading to data leakage. To mitigate this, we collect a new C/C++ function-based VD dataset from NVD [29], encompassing various vulnerability types and projects, with all samples having commit dates after the pre-training cutoff dates of PLMs evaluated in this study.
@@
To ensure the comprehensiveness and representativeness of the evaluation, we evaluate 17 PLMs with parameter sizes ranging from millions to billions, covering a variety of model architectures. All models have been specifically pre-trained on code structure-aware tasks or exposed to large-scale code corpora, making their evaluation on VD tasks particularly relevant and competitive. Standing out from other existing evaluation efforts, we comprehensively compare PLMs’ performance using two model adaptation techniques: fine-tuning and prompt engineering. Specifically, we fully fine-tune small language models (i.e., BERT series and CodeT5), while applying LoRA for partial fine-tuning of LLMs with up to 34B parameters. For prompting engineering, we adopt both zero-shot and few-shot prompting. Within each prompt setting, in addition to raw code functions, we introduce three types of structural and semantic-aware prompts: flattened abstract syntax tree, code with API calls, and code with data flow. These prompts embed structural information and dependencies within the code, guiding PLMs to better analyze vulnerabilities. Further, we evaluate the performance of PLMs on out-of-distribution data and test data under various perturbations (i.e., normalization, abstraction, semantic-preserving transformations) to examine their generalizability and robustness in practical applications.
@@
Through experimental analysis, we reveal the following findings: (1) PLMs that have been pre-trained on specialized tasks that guide them in learning code syntactic and semantic features (e.g., PDBERT [30]) significantly outperform most PLMs that have been pre-trained or fine-tuned on large code corpora (e.g., CodeLlama [31]), despite the former having far fewer parameters. This suggests that future research on PLMs for VD may strike a balanced between cost and performance. (2) Evaluating fine-tuned PLMs on test data derived from the same source as the training data may result in inaccurate assessments of the model capabilities for VD in real-world scenarios. (3) Existing PLMs continue to struggle with detecting vulnerabilities that involve complex program dependencies. (4) PLMs still lack robustness to minor perturbations, such as inconsistencies in normalization rules applied during training and testing. (5) Most PLMs demonstrate certain robustness to abstracted code, suggesting that their predictions do not rely solely on textual words. (6) Most PLMs exhibit varing degrees of performance drop in semantic-preserving transformations, indicating that they are not yet reliable against vulnerable code reuse or adversarial examples. (7) The limited context window size of PLMs unintentionally introduces label errors during truncation, disrupting model training. Code slicing that reduces input length can help PLMs focus more effectively on learning vulnerability patterns.
@@
  • We conduct extensive evaluations of VD capabilities of 17 representative PLMs, covering various architectures, parameter scales, and model adaptation techniques including fine-tuning (up to 34B parameters) and prompt engineering (2 settings × 4 types) on a reconstructed dataset with high-quality labeling and a newly collected dataset from NVD encompassing diverse vulnerability types and projects.
  • We examine the robustness of existing PLMs in real-world VD scenarios by applying code normalization, code abstraction, semantic-preserving transformation. providing a series of valuable insights for future research.
  • We implement a new framework to assess above capabilities of PLMs, and automatically generate leakage-free VD datasets for evaluating future models trained on more recent data. Our artifacts are available at RevisitVD.
TABLE I: Comparison of vulnerability detection benchmarks
@@
Adaptation Method Model Diversity Data Diversity Time Split Balanced Training Realistic Testing Knowledge Cutoff Finetuned Model Size Out of Distribution Robust Analysis
Khare [32] Prompt Engineering - - - - -
Steenhoek [19] Prompt Engineering - - - - -
SecLLMHolmes [20] Prompt Engineering - - - - -
CORRECT [25] Prompt Engineering - - - - -
VulnSage [26] Prompt Engineering - - - - -
VulnLLMEval [27] Prompt Engineering - - - - -
VulDetectBench [28] Prompt Engineering - - - - -
VulBench [21] Prompt Engineering - - - - -
Steenhoek [22] Prompt Engineering - - - - -
LLM4Vuln [23] Prompt Engineering - - - - -
SecureFalcon [24] Prompt Engineering - - - - -
Zhang [18] Prompt Engineering - - - - -
DiverseVul [1] Fine-tuning S
Thapa [2] Fine-tuning M
CleanVul [3] Fine-tuning M
VulLLM [4] Fine-tuning L
VulnPatchPairs [5] Fine-tuning S
Aleksei [6] Fine-tuning L
Steenhoek [7] Fine-tuning S
Jiang [8] Fine-tuning M
Ni [9] Both S
PrimeVul [33] Both XL
Yin [10] Both M
Zhang [16] Both M
VulEval [17] Both S
Purba [11] Both L
Zhou [12] Both M
Zhou [13] Both S
ChatGPT4Vul [14] Both S
Guo [15] Both M
RivisitVD (Ours) Both XL
  • S: 125-220M; M: 7-8B; L: 13-15B; XL: 33-34B
@@
II Background and Related Work
II-A Pre-trained Language Models for Code
The Transformer architecture [34] has become the foundation of most pre-trained language models, using self-attention to capture dependencies across entire sequences. It consists of an encoder and a decoder, which can be used independently or together depending on the task. For encoder-based PLMs, one of the core pre-training tasks is masked language modeling [35] where the model improves contextual understanding by learning to predict randomly masked tokens in the sequence. For decoder-based PLMs, the primary objective during pre-training is causal language modeling [36] where the model improves its understanding of language patterns by learning to predict the next token based on preceding context.
@@ II-A Pre-trained Language Models for Code
Given the unique syntactic and structural characteristics of code, PLMs originally designed for natural language processing (NLP) tasks often require specialized adaptations when applied to code-related tasks. These adaptations include: (1) introducing specifically designed code structure-aware pre-training tasks to enhance code understanding, and (2) pre-training PLMs on large code corpora that encompass source code from multiple programming languages. The former approach is typically implemented using encoder-based model architectures, with parameter sizes generally in the millions, while the latter often utilizes decoder-based model architectures, with parameter sizes reaching billions. In order to distinguish them, this paper refers to code-specific small PLMs with millions of parameters as Code SLMs and those with billions of parameters as LLMs. Additionally, LLMs can be further categorized as Code LLMs or general-purpose LLMs, depending on their intended application.
@@
Small Code Language Models (Code SLMs). Code LMs often integrate specialized, code-specific knowledge into their pre-training tasks. For instance, UniXCoder [37] transforms the Abstract Syntax Tree (AST) of a function into a flattened AST using its proposed mapping algorithm, aiding the model in learning syntactic information within code. GraphCodeBERT [38] integrates a data def-use prediction task during pre-training, enhancing the model’s capability to capture data flow. PDBERT [30] introduces two pre-training tasks: statement-level control dependency prediction and token-level data dependency prediction, to guide the model in learning semantic relationships in code. CodeT5 [39] involves an identifier-aware pre-training task to differentiate and recover identifiers for improved code understanding and generation.
@@
Large Code Language Models (Code LLMs). Within state-of-the-art Code LLMs, the fill-in-the-middle approach [40] is adopted by models like CodeLlama [31] and DeepSeek-Coder [41] to complement left-to-right generative capabilities of decoder-based PLMs, enhancing code generation and infilling. To handle long-context tasks such as cross-file code completion, many Code LLMs extend input sequence length by reconfiguring parameters in rotary position embeddings (RoPE) [42]. To further improve Code LLMs’ ability for code-related tasks, numerous curated instruction datasets are used to fine-tune Code LLMs [43, 44].
II-B Evaluating PLMs for Vulnerability Detection

### v3

@@
Abstract.
The rapid advancement of pre-trained language models (PLMs) has demonstrated promising results for various code-related tasks. However, their effectiveness in detecting real-world vulnerabilities remains a critical challenge. While existing empirical studies evaluate PLMs for vulnerability detection (VD), they suffer from data leakage, limited scope, and superficial analysis, hindering the accuracy and comprehensiveness of evaluations. This paper begins by revisiting the common issues in existing research on PLMs for VD through the evaluation pipeline. It then proceeds with an accurate and extensive evaluation of 18 PLMs, spanning model parameters from millions to billions, on high-quality datasets that feature accurate labeling, diverse vulnerability types, and various projects. Specifically, we compare the performance of PLMs under both fine-tuning and prompt engineering, assess their effectiveness and generalizability across various training and testing settings, and analyze their robustness to perturbations such as code normalization, abstraction, and semantic-preserving transformations.
Our findings reveal that, for function-level VD, PLMs incorporating pre-training tasks designed to capture the syntactic and semantic patterns of code outperform both general-purpose PLMs and those solely pre-trained or fine-tuned on large code corpora. However, these models face notable challenges in real-world scenarios, such as difficulties in detecting vulnerabilities with complex dependencies, handling perturbations introduced by code normalization and abstraction, and identifying semantic-preserving vulnerable code transformations. Also, the truncation caused by the limited context windows of PLMs can lead to a non-negligible number of labeling errors, which is overlooked by previous work. This study underscores the importance of thorough evaluations of model performance in practical scenarios and outlines future directions to help enhance the effectiveness of PLMs for realistic VD applications.
large language model, vulnerability detection, software security
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 21st ACM ASIA Conference on Computer and Communications Security; June 01–06, 2025; Bangalore, Indiaisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application security
1. Introduction
@@
Although existing research has explored if PLMs can demonstrate strong performance in vulnerability detection (VD), limitations across various stages of the evaluation pipeline hinders an accurate reflection of PLMs’ true capabilities: (i) Data Leakage. Most studies rely on evaluation datasets that inherently introduce data leakage, leading to biased estimations of the model performance in real-world scenarios. The randomly partitioning method used in existing studies (diversevul; a16; a17; a18; a19; a20; benjamin2023; a22; ni1; yinxin; a28; a29; davidlo; a31; a32) often causes models to encounter duplicated code patterns between the training and test data. The temporal overlap between the pre-training data cut-off dates of PLMs and the commit dates of evaluation data is also frequently overlooked (a26; a27; a28; a29; davidlo; a31; a32). (ii) Limited Scope. The experimental setup and settings adopted by many studies are neither comprehensive nor aligned with real-world scenarios, resulting in unrepresentative conclusions. For example, some studies focus on models with constrained architectures and parameter scales (graph_prompt; diversevul; a16; a17). Also, models are often evaluated on samples with limited size, a narrow range of vulnerability types (benjamin; ullah; a10; a11; a12; a13; a28; a29; davidlo), or an unrealistically balanced distribution (a6; a7; a8; a9; a10; a11; a12; a13). Others focus exclusively on either fine-tuning (diversevul; a16; a17; a18; a19; a20; benjamin2023; a22) or prompting engineering (a6; a7; a8; a9; a10; a11; a12; a13), neglecting comparative analyses across different adaptation methods or considerations of the trade-offs between cost and performance of PLMs in VD. (iii) Superficial Analysis. Existing studies (yinxin; a26; a27; a28; a29; davidlo; a31; a32) primarily focus on performance comparisons, without thoroughly investigating how practical factors (e.g., code normalization, abstraction, transformation, truncation) influence the effectiveness of PLMs for VD. This creates a significant gap between performance estimates and real-world applications, failing to provide insights for enhancing PLMs’ true capabilities.
@@
Therefore, this paper Revisits the capabilities of PLMs for VD (RevisitVD) through an extensive and realistic evaluation that addresses the shortcomings of existing evaluation works. In particular, starting from data preparation, we discuss the limitations of existing work in selecting evaluation datasets, including insufficient consideration of data volume, the diversity of vulnerability types, and inherent labeling errors within the datasets. To address these issues, we introduce our reconstructed dataset, alongside a time-order-based dataset partitioning method to avoid data leakage caused by the random data partitioning commonly used in current VD research. Additionally, we highlight the risk that existing VD datasets may predate the cutoff date of PLMs’ pre-training data, potentially leading to data leakage. To mitigate this, we collect a new C/C++ function-based VD dataset from NVD (nvd), encompassing various vulnerability types and projects, with all samples having commit dates after the pre-training cutoff dates of PLMs evaluated.
@@
To ensure the comprehensiveness and representativeness of the evaluation, we evaluate 18 PLMs with parameter sizes ranging from millions to billions, covering a variety of model architectures. All models have been specifically pre-trained on code structure-aware tasks or exposed to large-scale code corpora, making their evaluation on VD tasks particularly relevant and competitive. Standing out from other existing evaluation efforts, we comprehensively compare PLMs’ performance using two model adaptation techniques: fine-tuning and prompt engineering. Specifically, we fully fine-tune small language models (i.e., BERT series and CodeT5), while applying LoRA for partial fine-tuning of LLMs with up to 34B parameters. For prompting engineering, we adopt both zero-shot and few-shot prompting. Within each prompt setting, in addition to raw code functions, we introduce three types of structural and semantic-aware prompts: flattened abstract syntax tree, code with API calls, and code with data flow. These prompts embed structural information and dependencies within the code, guiding PLMs to better analyze vulnerabilities. We also introduce the existing state-of-the-art (SOTA) reasoning model, GPT-5 Mini, to explore the improvements that Chain-of-Thoughts (CoT) brings to VD. Furthermore, we evaluate the performance of PLMs on out-of-distribution data and test data under various perturbations (i.e., normalization, abstraction, semantic-preserving transformations) to examine their generalizability and robustness in practical applications.
@@
Through experimental analysis, RevisitVD reveals the following findings: (1) PLMs pre-trained on specialized tasks that guide them in learning code syntax and semantics (e.g., PDBERT) significantly outperform most PLMs pre-trained or fine-tuned on general language modeling (e.g., CodeLlama), despite having far fewer parameters. This suggests that future research could focus on designing specialized pre-training tasks to improve PLM capabilities for VD while considering the trade-off between performance and efficiency. (2) Evaluating fine-tuned PLMs on test data from the same source as the training data may lead to inaccurate assessments of their real-world capabilities. (3) Existing PLMs struggle with detecting vulnerabilities that involve complex program dependencies, and lack robustness to minor perturbations, such as inconsistencies in normalization rules applied during training and testing. (4) Most PLMs demonstrate certain robustness to abstracted code, suggesting that their predictions do not rely solely on textual words. (5) Most PLMs exhibit varing degrees of performance drop in semantic-preserving transformations, indicating that they are not yet reliable against vulnerable code reuse or adversarial examples. (6) The limited context window size of PLMs unintentionally introduces label errors during truncation, disrupting model training. Code slicing that reduces input length can help PLMs focus more effectively on learning vulnerability patterns.
@@
(1) We revisit the true capabilities of PLMs for VD by addressing multiple critical shortcomings of existing research. These include data leakage caused by flawed dataset partitioning and overlooking the temporal overlap between evaluation data and pre-training knowledge of LLMs; limited scope resulting from the evaluation of constrained evaluation setups on unrepresentative experimental settings; and superficial analysis neglecting factors like model generalization and robustness to practical use.
(2) We evaluate a wide range of 18 PLMs, from small to large parameter scales, and compares their performance using two adaptation techniques: fine-tuning and prompt engineering. These comprehensive comparisons, which also includes the use of structural and semantic-aware prompts and the CoT reasoning model, offer a realistic estimation of PLMs’ capabilities for VD.
(3) We implement an extensible evaluation framework built on newly collected, high-quality datasets that go beyond LLMs’ pre-training knowledge cutoff, allowing for future evaluation of stronger models on more recent data. We reveal several key insights into the capabilities and weaknesses of PLMs, along with effective strategies for enhancing their performance in practical settings. Our artifacts are available https://github.com/youpengl/RevisitVD.
2. Background and Related Work
2.1. Pre-trained Language Models
The Transformer architecture (attention) has become the foundation of most pre-trained language models, using self-attention to capture dependencies across entire sequences. It consists of an encoder and a decoder, which can be used independently or together depending on the task.
Encoder-based Pre-trained Language Models. For encoder-based PLMs, one of the core pre-training tasks is masked language modeling (bert) where the model improves contextual understanding by learning to predict randomly masked tokens in the sequence. Given a token sequence $\mathcal{U}^{mask}=\{u_{0},u_{1}^{mask},\cdots,u_{i}^{mask},\cdots,u_{m-1}\}$ with a certain percentage of randomly masked tokens, the optimization of the PLM $\Theta$ can be formulated as follows:
(1) $\Theta=\arg\max_{\Theta}\sum_{u_{i}\in S_{mask}}\log P(u_{i}\mid\mathcal{U}^{mask};\Theta),$
where $S_{mask}$ is the set of masked tokens and $\Theta$ represents the parameters of the PLM, optimized through stochastic gradient descent (SGD).
Decoder-based Pre-trained Language Models. For decoder-based PLMs, the primary objective during pre-training is causal language modeling (GPT) where the model improves its understanding of language patterns by learning to predict the next token based on preceding context. Given a sequence of tokens $\mathcal{U}=\{u_{0},u_{1},\cdots,u_{i},\cdots,u_{m-1}\}$, the optimization of the PLM $\Theta$ can be expressed as follows:
(2) $\Theta=\arg\max_{\Theta}\sum_{i}\log P(u_{i}\mid u_{i-n},\ldots,u_{i-1};\Theta),$
where $n$ denotes the size of the context window, and $\Theta$ represents the parameters of the PLM.
Table 1. Comparison of vulnerability detection benchmarks
@@ 2.1. Pre-trained Language Models
Model Diversity Data Diversity Time Split Balanced Training Realistic Testing Knowledge Cutoff Finetuned Model Size Out of Distribution Robust Analysis
Prompt Engineering-based Evaluation
Khare et al (understanding) N/A N/A N/A N/A N/A
Steenhoek et al (benjamin) N/A N/A N/A N/A N/A
SecLLMHolmes (ullah) N/A N/A N/A N/A N/A
CORRECT (a6) N/A N/A N/A N/A N/A
VulnSage (a7) N/A N/A N/A N/A N/A
VulnLLMEval (a8) N/A N/A N/A N/A N/A
VulDetectBench (a9) N/A N/A N/A N/A N/A
VulBench (a10) N/A N/A N/A N/A N/A
Steenhoek et al (a11) N/A N/A N/A N/A N/A
LLM4Vuln (a12) N/A N/A N/A N/A N/A
SecureFalcon (a13) N/A N/A N/A N/A N/A
Zhang et al (graph_prompt) N/A N/A N/A N/A N/A
Fine Tuning-based Evaluation
DiverseVul (diversevul) $\sim$ 220M
Thapa et al (a16) $\sim$ 6B
CleanVul (a17) $\sim$ 7B
VulLLM (a18) $\sim$ 15B
VulnPatchPairs (a19) $\sim$ 220M
Aleksei et al (a20) $\sim$ 13B
Steenhoek et al (benjamin2023) $\sim$ 125M
Jiang et al (a22) $\sim$ 8B
Considering Both Prompt Engineering- and Fine Tuning-based Evaluation
Ni et al (ni1) $\sim$ 125M
PrimeVul (primevul) $\sim$ 175B
Yin et al (yinxin) $\sim$ 7B
Zhang et al (a26) $\sim$ 8.5B
VulEval (a27) $\sim$ 220M
Purba et al (a28) $\sim$ 13B
Zhou et al (a29) $\sim$ 8B
Zhou et al (davidlo) $\sim$ 125M
ChatGPT4Vul (a31) $\sim$ 125M
Guo et al (a32) $\sim$ 7B
RevisitVD (Ours) $\sim$ 34B
@@
2.2. Pre-trained Language Models for Code
@@
Given the unique syntactic and structural characteristics of code, PLMs originally designed for natural language processing (NLP) tasks often require specialized adaptations when applied to code-related tasks. These adaptations include: (1) introducing specifically designed code structure-aware pre-training tasks to enhance code understanding, and (2) pre-training PLMs with the general language modeling objective on large code corpora that encompass source code from multiple programming languages. The former approach is typically implemented using encoder-based model architectures, with parameter sizes generally in the millions, while the latter often utilizes decoder-based model architectures, with parameter sizes reaching billions. In order to distinguish them, this paper refers to code-specific small PLMs with millions of parameters as Code SLMs and those with billions of parameters as LLMs. Additionally, LLMs can be further categorized as Code LLMs or general-purpose LLMs, depending on their intended application.
@@
Small Language Models for Code (Code SLMs). Code SLMs often integrate specialized, code-specific knowledge into their pre-training tasks. For instance, UniXCoder (UniXcoder) transforms the Abstract Syntax Tree (AST) of a function into a flattened AST using its proposed mapping algorithm, aiding the model in learning syntactic information within code. GraphCodeBERT (graphcodebert) integrates a data def-use prediction task during pre-training, enhancing the model’s capability to capture data flow. PDBERT (pdbert) introduces two pre-training tasks: statement-level control dependency prediction and token-level data dependency prediction, to guide the model in learning semantic relationships in code. CodeT5 (codet5) involves an identifier-aware pre-training task to differentiate and recover identifiers for improved code understanding and generation.
@@
Large Language Models for Code (Code LLMs). For Code LLMs, the fill-in-the-middle approach (FIIM) is adopted by models like CodeLlama (codellama) and DeepSeek-Coder (deepseekcoder) to complement left-to-right generative capabilities of decoder-based PLMs, enhancing code generation and infilling. To handle long-context tasks such as cross-file code completion, many Code LLMs extend input sequence length by reconfiguring parameters in rotary position embeddings (RoPE) (ROPE). To further improve Code LLMs’ ability for code-related tasks, numerous curated instruction datasets are used to fine-tune Code LLMs (starchat; wizardcoder).
2.3. Evaluating PLMs for Vulnerability Detection